answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

How many arrangements of the letters MATH are possible if each letter can be used only once in each arrangement?

24

The first letter can be any of the 4. The second can be any of the remaining 3. The third, either remaining, and the last, the only remaining one. This means there are 4*3*2*1=24 different arrangements.

If the standard deviation decreases does the mean decrease?

Not necessarily. The standard deviation measures (in simplified terms) how different the numbers are from each other, while the mean is their average. If the standard deviation decreases, it means the numbers are closer to each other, it doesn't change how big the numbers are.

What is the sample space of a coin tossed n times?

The sample space consists of 2n ordered n-tuples of the form (X1, X2, ..., Xn) where each Xi = H or T.

What is obtained by squaring the standard deviation?

Variance. Although usually it is the SD that is calculated from the variance.

What is the probability that the event will not happen if the probability of an event is 75 percent?

let event is X

so

P(X)=75%=0.75

probabilty that event will not happen

P(X')=1-P(X)

P(X')=1-0.75

P(X')=0.25

so

probabilty that event will not happen is 0.25

A coin is tossed 10 times what is the probability of getting all tails?

The answer is 1 in 2^10 or 1 in 1024 or approximately 0.098%.

While there is a 50% chance of getting a heads or tails in a FAIR coin toss during each toss, there is a very low likelihood of getting 10 in row as follows:

1/2 * 1/2 * 1/2 * 1/2 * 1/2 * 1/2 * 1/2 * 1/2 * 1/2 * 1/2 = 1 / 1024 = approximately 0.098%.

Can average deviation be negative?

No, average deviation cannot be negative.

Deviation is a representation of differences between numbers. A difference is always an absolute value, so the number cannot be negative (even though subtracting the deviation from an average may result in a a negative result).

What is converse probability?

It's deduction. You can find the odds of something happening by first finding out the odds of that something not happening. That is converse probability. For instance the odds of rolling a "3" on a 6 sided die. Using converse probability would be 5/6 (5 sides are not the number "3"). 6/6 (six sides in all) - 5/6 = 1/6 is the odds of rolling that "3".

What season are rapes more likely to occur?

It has been confirmed that in the summer time more rapes happen. This is because during the summer kids are on their summer vacations so therefore means twice the rapes.

What is the chance of drawing a red ball from a bag containing 3 red and 2 white balls?

The probability is 1 if you draw three balls without replacement.

If only one draw, it is 3/5.

How would you separate a red and blue marbles of the same size and mass?

At the simplest level, one would analyze the color of each marble, creating two piles of marbles (one red and one blue).

If the person was not able to see colors, they might consider using a spectrometer (spectrograph). Measuring each marble individually, each marble would be placed in the device and the wavelength would be displayed by the device. Based on the wavelength (red is in the range of 620-750nm and blue is in the range of 450-475nm), the person would still create two piles of marbles, the higher wavelength one (indicating red) and the lower wavelength one (indicating blue).

How many possible outcomes would there be if four coins were tossed once?

there could be five possible outcomes: t= tails h= heads

t t t t 4 tails

t t t h 3 tails 1 head

t t h h 2 tails 2 heads

t h h h 1 tail 3 heads

h h h h 4 heads

To pass a true or false test a student must answer 8 out of 10 correct. What is the probability a student will pass if he quesses on all 10 questions?

Total number of different ways he can fill out the answer sheet = 210 = 1,024 .

He passes the test if he gets zero wrong, one wrong, or two wrong.

Number of ways to get zero wrong = 1.

Number of ways to get exactly one wrong = 10.

Number of ways to get exactly two wrong = 45.

Total number of ways to get exactly 0, 1, or 2 wrong = (1 + 10 + 45) = 56.

Probability of passing = (56) / (1,024) = 5.47%(rounded)

Is it possible to roll a die five times and get a six every time?

Yes is it possible but not probable

note: please dont edit this question with stupid awnsers kthxbye

How is the expected value of a normal distribution computed?

If a random variable X is distributed normally with probability distribution function p(x), then the expected value of X is

E(X) = integral of x*p(x)dx evaluated over the whole of the real line.

Why is pelton wheel not suitable for low heads?

low head will generate low velocity (since velocity is the function of head) which results into the less discharge.

therefore the power generated will be less

hence is not suitable

What is an acceptable standard deviation?

An acceptable standard deviation depends entirely on the study and person asking for the study. The smaller the standard deviation, the more acceptable it will be because the less likely there are to be errors.